OpenStudy (anonymous):

Can someone please help me with this? How do I write this expression in simplified radical form?

OpenStudy (anonymous):

\[\frac{ 3-2\sqrt{11} }{ 2+\sqrt{11} }\]

OpenStudy (anonymous):

Can someone just tell me where to start? I honestly have no clue.